my question is A Rectangular Persian carpet has a perimeter of 240 inches. The length of the carpet is 25 inches more then the width. What are the dimensions of the carpet? Solve algebraically
Answer by ikleyn(52781)

.

As you, probably, know, the perimeter is the sum of a rectangle side lengths


    Perimeter = L + W + L + W = 2L + 2W.


In this problem, we are given  L = W + 25  inches.


Therefore,


    Perimeter = 2*(W+25) + 2W = 240  inches.


Simplify and solve for W, which is the width of your Rectangular Persian carpet.


    2W + 50 + 2W = 240

    4W           = 240-50 = 190

     W                    = 190/4 = 47.5 inches.


Thus the width of your Rectangular Persian carpet is  47.5 inches.


Hence, the length of your Rectangular Persian carpet is 25 inches more,  47.5 + 25 = 72.5.


ANSWER.  The dimensions of your Rectangular Persian carpet are  72.5 inches (the length) and 47.5 inches (the width).


CHECK:    72.5
          47.5
          72.5
          47.5
   -----------------

         240.0 inches.      ! Precisely correct !

Solved.